# Thermodynamic theory of square skyrmion lattice in tetragonal frustrated antiferromagnets

### v1

@@ Abstract
High-temperature part of the phase diagram of tetragonal frustrated antiferromagnets is discussed within the framework of the mean-field approach. Based on recent experimental and theoretical studies, we consider a model including frustrated Heisenberg exchange, biquadratic exchange, anisotropic exchange, and magnetodipolar interaction. It is analytically demonstrated that a subtle interplay among these interactions results in a variety of phase diagrams in the temperature-magnetic field plane. We argue that the phase diagram obtained for the easy axis along the high-symmetry direction $[001]$ and moderate biquadratic exchange reproduces all crucial features of the phase diagram experimentally observed for GdRu2Si2 compound. The topological properties of the respective double-modulated square skyrmion lattice phase are discussed. The developed analytical approach can also be useful for the refinement of the microscopic model parameters when comparing its predictions with experimental findings.
keywords:
Square skyrmion lattice , Topological phases , Phase diagram , Mean-field approach , Biquadratic exchange
journal: Annals of Physics
[…]
[first]organization=Center for Theoretical Physics of Complex Systems, Institute for Basic Science, city=Daejeon, postcode=34126, country=Republic of Korea
[…]
[second]organization=Department of Physics, Saint-Petersburg State University, city=St. Petersburg, postcode=198504, country=Russian Federation
1 Introduction
@@
Skyrmions and their ordered arrays – skyrmion lattices (SkL) – are among the most studied objects of contemporary condensed matter physics Bogdanov and Panagopoulos (2020). Skyrmions were first predicted as metastable states of 2D ferromagnets Belavin and Polyakov (1975). A possibility to stabilize isolated skyrmions due to Lifshitz invariants [Dzyaloshinskii-Moriya interaction Dzyaloshinsky (1958); Moriya (1960)] was shown in Ref. Bogdanov and Yablonskii (1989). Next, it was demonstrated that skyrmions can form an ordered state, a hexagonal skyrmion lattice, under certain conditions Bogdanov and Hubert (1994). Notably, this superstructure was indeed observed in MnSi using neutron scattering Mühlbauer et al. (2009) and the Hall effect Neubauer et al. (2009), which caused an upsurge of interest in topological spin structures. It is also stimulated by prospective applications, e.g., the racetrack memory Fert et al. (2013, 2017) or unconventional computing Li et al. (2021).
@@
The efficiency of various applications relies on magnetic skyrmions’ non-trivial topology Belavin and Polyakov (1975) and, hence, topological protection from local perturbations. The topological charge of a spin structure is defined using the spin winding number on a unit sphere:
[…]
@@
where $\mathbf{n}=\mathbf{s}/|\mathbf{s}|$ is a unit vector along the averaged over thermodynamical and quantum fluctuations spin direction. For a single skyrmion, the integral usually yields $Q=\pm 1$, whereas for SkL the natural measure is a density of the topological charge, $n_{\textrm{sk}}$. The latter quantity is crucial. For instance, it defines the topological Hall resistivity ($\rho^{T}_{xy}\propto n_{\textrm{sk}}$eubauer et al. (2009). Nowadays, other topologically nontrivial objects, e.g., meron-antimeron structures, are also actively studied Yu et al. (2018); Göbel et al. (2021); Hayami and Yambe (2021); Kim et al. (2023); Kuchkin et al. (2023); Leonov (2024).
@@
Remarkably, frustration of exchange interaction can lead to the stabilization of short-period magnetic structures with a high density of topological charge $n_{\textrm{sk}}$ Leonov and Mostovoy (2015), leading to the giant topological Hall effect. Compact skyrmion lattices with periods of several nanometers were indeed observed in triangular Gd2PdSi3 Kurumaji et al. (2019), breathing kagomé Gd3Ru4Al12 Hirschberger et al. (2019), and tetragonal GdRu2Si2 Khanh et al. (2020), GdRu2Ge2 Yoshimochi et al. (2024), EuAl4 Takagi et al. (2022). Recent experimental and theoretical progress on this topic is reviewed in Ref. Kawamura (2025).
@@
Shortly after the square skyrmion lattice was observed in GdRu2Si2 [see Ref. Khanh et al. (2020)], three theoretical studies addressing mechanisms of its stabilization were published Hayami and Motome (2021); Utesov (2021); Wang et al. (2021). It was shown that important ingredients of the corresponding microscopic models are biquadratic exchange Takahashi (1977); Spišák and Hafner (1997); Hayami et al. (2017) and complex anisotropic interactions. The latter include compass anisotropy, anisotropic exchange, and magnetodipolar interaction. The subsequent studies were devoted to the identification of various magnetic phases Khanh et al. (2022); Wood et al. (2023), electronic structure, and exchange couplings Bouaziz et al. (2022); Wood et al. (2025) of GdRu2Si2.
@@
In the present study, based on recent experimental findings, we revisit the high-temperature part of the phase diagram of tetragonal frustrated antiferromagnets. We upgrade the mean-field approach of Ref. Utesov (2021) to include the biquadratic exchange interaction. We show that its interplay with complex magnetic anisotropies results in a plethora of phase diagrams in the temperature–magnetic field plane. Using the analytical expressions for various spin structures’ free energies, we discuss their competition. It is shown that the existence of higher-order harmonics is crucial for the stabilization of the square SkL due to a peculiar correction to the free energy, which resembles the corresponding term for hexagonal SkLs arising in external field due to the condition $\mathbf{k}_{1}+\mathbf{k}_{2}+\mathbf{k}_{3}=0$ [see, e.g., Refs. Mühlbauer et al. (2009); Leonov and Mostovoy (2015); Utesov (2022)]. In our case, we denote modulation vectors of two principal helical constituents of SkL as $\mathbf{k}_{x,y}$ and the higher harmonics ones as $\mathbf{k}_{\pm}=\mathbf{k}_{x}\pm\mathbf{k}_{y}$. The corresponding condition reads $\mathbf{k}_{x}\pm\mathbf{k}_{y}-\mathbf{k}_{\pm}=0$. The same conclusion was also emphasized in the recent numerical study Hayami (2025). Exploring the parameter space, we find that in the case of easy axis along the high-symmetry tetragonal axis and in a certain range of biquadratic exchange constant, the phase diagrams similar to the experimental one of GdRu2Si2 Garnier et al. (1996); Khanh et al. (2020, 2022) can be obtained.
@@
The rest of the paper is organized as follows. In Section 2, we introduce the formalism. Various possible spin structures and relations between their free energies are discussed in Sec. 3. In Sec. 4, we present particular phase diagrams, including the one corresponding to GdRu2Si2. Section 5 is devoted to a general discussion of the results. Section 6 contains our conclusions and summary. Finally, two appendices contain detailed information about free energy calculations and topological properties of some spin structures.
2 General formulae
@@
Following the previous studies Hayami and Motome (2021); Utesov (2021); Wang et al. (2021), we consider a model which includes frustrated isotropic exchange, magnetodipolar interaction and/or anisotropic exchange, Zeeman term, and biquadratic exchange. Assuming one magnetic ion in the unit cell, we can write the reciprocal space Hamiltonian as follows:
@@
where $N$ is the number of lattice sites and we measure magnetic fields in energy units, $h=g\mu_{\textrm{B}}H$. The last term with $K>0$ here is written in the form suggested in Refs. Hayami et al. (2017); Hayami and Motome (2021) as the most important contribution to the biquadratic exchange in itinerant magnets. The first two terms (3) and (4) can be combined into a single term $\mathcal{H}_{\textrm{EX}}+\mathcal{H}_{\textrm{AN}}=-\sum_{\mathbf{q}}\mathcal{H}^{\alpha\beta}_{\mathbf{q}}S^{\alpha}_{\mathbf{q}}S^{\beta}_{-\mathbf{q}}$. Noteworthy, tensor $\mathcal{H}^{\alpha\beta}_{\mathbf{q}}$ eigenvalues $\lambda_{1}(\mathbf{q})\geq\lambda_{2}(\mathbf{q})\geq\lambda_{3}(\mathbf{q})$ and respective eigenvectors $\mathbf{v}_{1}(\mathbf{q}),\,\mathbf{v}_{2}(\mathbf{q}),\,\mathbf{v}_{3}(\mathbf{q})$ define the “zoo” of possible spin structures in each particular model.
@@
We aim to discuss experimentally relevant cases, when the possible modulation vectors lie in-plane [we denote them as $\mathbf{k}_{x}=(k,0,0)$ and $\mathbf{k}_{y}=(0,k,0)$]. They are equivalent due to tetragonal symmetry Khanh et al. (2020, 2022). These vectors are roughly (up to small anisotropy-induced corrections) determined by the maxima of the exchange interaction Fourier transform. In the discussion below, we will also need $\mathbf{k}_{\pm}=\mathbf{k}_{x}\pm\mathbf{k}_{y}$. In GdRu2Si2, $k=0.22\times 2\pi/a$, where $a$ is the in-plane lattice parameter Khanh et al. (2020). Moreover, the corresponding spin structures have transverse order parameters, and we conclude that the longitudinal directions for $\mathbf{k}_{x}$ and $\mathbf{k}_{y}$ are the respective hard axes. Then, the corresponding perpendicular in-plane directions can be easy or middle axes. Importantly, the former case is relevant to magnetodipolar interaction, which should be important for Gd3+ ions with zero orbital momentum Utesov (2021). In the general case, the anisotropic part of the Hamiltonian can also include compass anisotropic exchange Wang et al. (2021) and single-ion terms. Fig. 1 illustrates these crucial concepts.
Figure 1: (a) Tetragonal crystalline lattice of GdRu2Si2  where only magnetic Gd3+ ions are shown. (b) According to the experimental observations Khanh et al. (2020, 2022), GdRu2Si2  has ordered magnetic phases with modulation vectors $\mathbf{k}_{x,y}$. Here we also show vectors $\mathbf{k}_{\pm}=\mathbf{k}_{x}\pm\mathbf{k}_{y}$. The respective harmonics are crucial for the square skyrmion lattice stabilization. (c) Momentum-dependent biaxial anisotropy (consisting of magnetodipolar, single-ion, and compass anisotropic exchange contributions) favors vertical screw spirals for $\mathbf{k}_{x,y}$ modulations.
@@
For our calculations, we employ the mean-field approach based on the smallness of the thermodynamically average ordered spin value, $s=\langle S\rangle\ll S$. In this case, which corresponds to high temperatures near $T_{\textrm{N}}$, we can express the free energy as a function of $s_{\mathbf{q}}$ (reciprocal space) and $\mathbf{s}_{i}$ (direct space) as follows [see, e.g., Ref. Gekht (1984) for details]:
$\mathcal{F}=-\sum_{\mathbf{q}}\mathcal{H}^{\alpha\beta}_{\mathbf{q}}s^{\alpha}_{\mathbf{q}}s^{\beta}_{-\mathbf{q}}-\sqrt{N}\mathbf{h}\cdot\mathbf{s}_{\bf 0}+AT\sum_{i}s^{2}_{i}+\frac{K}{N}\sum_{\mathbf{q}}(\mathbf{s}_{\mathbf{q}}\cdot\mathbf{s}_{-\mathbf{q}})^{2}+BT_{\textrm{N}}\sum_{i}s^{4}_{i},$ (7)
@@
For $S=7/2$, one has $A\approx 0.095$ and $B\approx 0.002$. Note that by assuming a certain spin structure, we readily obtain its free energy (7) in the form of a quartic polynomial suitable for either analytical or numerical minimization Gekht (1984); Utesov (2021, 2022).
3 Spin structures and phase transitions
In the present study, we consider the magnetic field only along the $\mathbf{c}$ axis. Some of the relevant spin structures are shown in Fig. 2. We distinguish: (i) single-modulated spin-density wave (SDW) which we denote 1S, it turns out to be relevant only for easy axes along $\hat{z}$, (ii) three variants of double-modulated superpositions of two SDWs with different polarizations (2SXY, 2SYZ, 2SZZ), (iii) vertical screw helicoid (1Q), (iv) transverse cone spiral (TC), (v) hybrid structure made of vertical spiral and SDW with perpendicular modulation vector (1Q+1S), (vi) superposition of two vertical spirals (2Q), which is the square skyrmion lattice in a certain range of fields, (vii) alternative 2Q ordering (2Q), where $\hat{z}$ components of the spin structure have $\mathbf{k}_{\pm}$ modulation vectors. Details of calculations supporting our claims are presented in A.
@@ 3 Spin structures and phase transitions
Figure 2: Sketch of main spin structures in the $xy$ plane, which can appear in typical phase diagrams of tetragonal frustrated antiferromagnets. Color encodes $\hat{z}$ projections of spins (red – “spin up”, violet – “spin down”, green – spin in the $xy$ plane). (a) Superposition of two spin-density waves with modulation vectors $\mathbf{k}_{x}$ and $\mathbf{k}_{y}$ and perpendicular in-plane polarizations (2SXY). In real space, it corresponds to a vortex lattice. (b) Superpositions of two SDWs with in-plane and out-of-plane polarizations (S2YZ). (c) Superposition of screw helicoid with $\mathbf{k}_{y}$ and SDW with $\mathbf{k}_{x}$ (1Q+1S); meron – anti-meron lattice. (d) Topologically-nontrivial square skyrmion lattice made of two screw helicoids with $\mathbf{k}_{x}$ and $\mathbf{k}_{y}$, harmonics with $\mathbf{k}_{\pm}$ polarized along $\hat{z}$, and uniform magnetization (2Q). Its topological charge per unit cell is $n_{\textrm{sk}}=1$ (e) Topologically-trivial version of 2Q structure at stronger external field (all $s^{z}_{i}>0$), $n_{\textrm{sk}}=0$. (f) Alternative topologically-nontrivial structure 2Q with $n_{\textrm{sk}}=2$.
@@
Moreover, it can be shown that in the region where 1Q+1S does not exist [$t<(5b+K)\Lambda_{1}/2b$], but 1Q does [$t>(3b+K)\Lambda_{1}/2b$], 2SXY has minimal free energy. As a corollary, the 1Q structure does not appear in the phase diagrams.
The same applies if we compare the free energies of 1Q+1S (51) and 2Q (55) [without the correction Eq. (59) important at moderate magnetic fields]. Note that these phases are defined in the same region $t>(5b+K)\Lambda_{1}/2b$ and emerge via the second-order phase transition from 2SXY. The difference in the free energies
$f_{\textrm{2Q}}-f_{\textrm{1Q+1S}}=\frac{(2bt-5b\Lambda_{1}-K\Lambda_{1})^{2}}{8(9b+2K)(2b^{2}+5bK+K^{2})}\geq 0.$ (11)
@@
$\mathbf{s}_{j}$ $=$ $s(\hat{y}\sin{\mathbf{k}_{x}\mathbf{R}_{j}}+\hat{x}\sin{\mathbf{k}_{y}\mathbf{R}_{j}})+\hat{z}[p(\cos{\mathbf{k}_{x}\mathbf{R}_{j}}+\cos{\mathbf{k}_{y}\mathbf{R}_{j}})+m],$ (12)
with certain $s,p$ and $m$ (see A). $z$-components of spin structure due to the quartic term in the free energy (7) lead to instability for the emergence of additional harmonics $\mathbf{k}_{\pm}$ lowering the total free energy. The physics here resembles the conventional mechanism of the stabilization of hexagonal skyrmion lattices in magnetic field due to interplay of constant magnetization and three helical constituents with $\mathbf{k}_{1}+\mathbf{k}_{2}+\mathbf{k}_{3}=0$ [see, e.g., Refs. Gekht (1984); Mühlbauer et al. (2009); Leonov and Mostovoy (2015); Utesov (2022)]. In the present case, the crucial terms $\propto mp^{2}$ emerge due to the equality $\mathbf{k}_{x}\pm\mathbf{k}_{y}-\mathbf{k}_{\pm}=0$ and similar ones. Notably, the corresponding correction to the spin structure
@@
“patches” the holes (where $\mathbf{s}_{j}=m\hat{z}$) in the modulated part of (12) for $\mathbf{k}_{x,y}\mathbf{R}_{j}=\pi l_{x,y}$ and $\mathbf{k}_{x}\mathbf{R}_{j}-\mathbf{k}_{y}\mathbf{R}_{j}=2\pi l+\pi$ (all $l$s are integers). Importance of these harmonics was also highlighted in Refs. Hayami (2022, 2025).
In the external field, the free energy parameters should be altered as follows: $t\rightarrow t^{\prime}=t-2b\chi^{2}h^{2}$ and $\Lambda_{1}\rightarrow\Lambda^{\prime}_{1}=\Lambda_{1}+4b\chi^{2}h^{2}$, where $\chi$ represents the uniform susceptibility. All free energies also acquire a “paramagnetic” correction $-\chi h^{2}/2$. So, in the external field, the above statements are intact. However, note that the correction to the 2Q phase free energy (59) is $\propto m^{2}$ and becomes crucial at moderate fields for which both $m$ and $p$ are significant.
@@
Importantly, near the transition point from 2SXY, for 2Q structure $p\ll 1$ (and, hence, $s_{\pm}\ll 1$), uniform magnetization $m$ dominates in $s^{z}_{i}$ and 2Q structure is topologically trivial, see B for details. The condition of nonzero topological charge is $2(p+s_{\pm})>m$. Note also that the transverse cone spiral, involving the hard magnetization axis, can be stabilized at lower $T$ [it is governed by the $\Lambda_{2}>\Lambda_{1}$ parameter, see Eq. (47)]. Its competition with the other phases should be checked separately. Particular examples of phase diagrams are presented below.
@@
4 Phase diagrams and application to GdRu2Si2

### v2

@@ Abstract
High-temperature part of the phase diagram of tetragonal frustrated antiferromagnets is discussed within the framework of the mean-field approach. Based on recent experimental findings, we generalize previous theoretical studies by considering a model that includes frustrated Heisenberg exchange, biquadratic exchange, magnetodipolar interaction, anisotropic exchange, and single-ion anisotropy. It is analytically demonstrated that a subtle interplay among these interactions results in a variety of phase diagrams in the temperature-magnetic field plane. We argue that one of the proposed diagrams reproduces all crucial features of the phase diagram experimentally observed for GdRu2Si2 compound. Besides magnetodipolar interaction and additional easy-axis contribution, it requires moderate biquadratic exchange. We show that despite the remarkable square skyrmion lattice being stable even if only magnetodipolar interaction/compass anisotropy or biquadratic exchange is included, their “symbiosis” allows for greatly enhancing its stability region. It is also demonstrated that higher-order harmonics play an important role in the stabilization of the square skyrmion lattice. The developed analytical approach can be useful for the refinement of the microscopic model parameters when comparing its predictions with experimental findings.
I Introduction
@@
Skyrmions and their ordered arrays – skyrmion lattices (SkL) – are among the most studied objects of contemporary condensed matter physics [1]. Skyrmions were first predicted as metastable states of 2D ferromagnets [2]. A possibility to stabilize isolated skyrmions due to Lifshitz invariants (Dzyaloshinskii-Moriya interaction [3, 4]) was shown in Ref. [5]. Next, it was demonstrated that skyrmions can form an ordered state, a hexagonal skyrmion lattice, under certain conditions [6]. Notably, this superstructure was indeed observed in MnSi using neutron scattering [7] and the Hall effect [8], which caused an upsurge of interest in topological spin structures. It is also stimulated by prospective applications, e.g., the racetrack memory [9, 10] or unconventional computing [11].
@@
The efficiency of various applications relies on magnetic skyrmions’ non-trivial topology [2] and, hence, topological protection from local perturbations. The topological charge of a spin structure is defined using the spin winding number on a unit sphere:
$Q=\frac{1}{4\pi}\int\mathbf{n}\cdot\left[\partial_{x}\mathbf{n}\times\partial_{y}\mathbf{n}\right]dxdy,$ (1)
@@
where $\mathbf{n}=\mathbf{s}/|\mathbf{s}|$ is a unit vector along the averaged over thermodynamical and quantum fluctuations spin direction. For a single skyrmion, the integral usually yields $Q=\pm 1$, whereas for SkL the natural measure is a density of the topological charge, $n_{\textrm{sk}}$. The latter quantity is crucial. For instance, it defines the topological Hall resistivity ($\rho^{T}_{xy}\propto n_{\textrm{sk}}$[8]. Nowadays, other topologically nontrivial objects, e.g., meron-antimeron structures, are also actively studied [12, 13, 14, 15, 16, 17].
@@
Remarkably, frustration of exchange interaction can lead to the stabilization of short-period magnetic structures with a high density of topological charge $n_{\textrm{sk}}$ [18], leading to the giant topological Hall effect. Compact skyrmion lattices with periods of several nanometers were indeed observed in triangular Gd2PdSi3 [19], breathing kagomé Gd3Ru4Al12 [20], and tetragonal GdRu2Si2 [21], GdRu2Ge2 [22], EuAl4 [23]. Recent experimental and theoretical progress on this topic is reviewed in Ref. [24].
@@
Shortly after the square skyrmion lattice was observed in GdRu2Si2 (see Ref. [21]), three theoretical studies addressing mechanisms of its stabilization were published [25, 26, 27]. It was shown that important ingredients of the corresponding microscopic models are biquadratic exchange [28, 29, 30] and complex anisotropic interactions. The latter include compass anisotropy, anisotropic exchange, and magnetodipolar interaction. The subsequent studies were devoted to the identification of various magnetic phases [31, 32], electronic structure, and exchange couplings [33, 34] of GdRu2Si2.
Figure 1: (a) Tetragonal crystalline lattice of GdRu2Si2  where only magnetic Gd3+ ions are shown. (b) According to the experimental observations [21, 31], GdRu2Si2  has ordered magnetic phases with modulation vectors $\mathbf{k}_{x,y}$. Here we also show vectors $\mathbf{k}_{\pm}=\mathbf{k}_{x}\pm\mathbf{k}_{y}$. The respective harmonics play an important role in the square skyrmion lattice stabilization. (c) Momentum-dependent biaxial anisotropy (consisting of magnetodipolar, single-ion, and compass anisotropic exchange contributions) favors vertical screw spirals for $\mathbf{k}_{x,y}$ modulations.
@@
In the present study, based on recent experimental findings, we revisit the high-temperature part of the phase diagram of tetragonal frustrated antiferromagnets. We generalize the mean-field approach of Ref. [26] to include the effects of the biquadratic exchange interaction into consideration. We show that its interplay with complex magnetic anisotropies results in a plethora of phase diagrams in the temperature–magnetic field plane. Using the analytical expressions for various spin structures’ free energies, we discuss their competition. It is shown that the existence of higher-order harmonics is crucial for the stabilization of the square SkL due to a peculiar correction to the free energy, which resembles the corresponding term for hexagonal SkLs arising in external field due to the condition $\mathbf{k}_{1}+\mathbf{k}_{2}+\mathbf{k}_{3}=0$ (see, e.g., Refs. [7, 18, 35]). In our case, we denote modulation vectors of two principal helical constituents of SkL as $\mathbf{k}_{x,y}$ and the higher harmonics ones as $\mathbf{k}_{\pm}=\mathbf{k}_{x}\pm\mathbf{k}_{y}$. The corresponding condition reads $\mathbf{k}_{x}\pm\mathbf{k}_{y}-\mathbf{k}_{\pm}=0$. The same conclusion was also emphasized in the recent numerical study [36]. Exploring the parameter space, we find that in the case of magnetodipolar interaction accompanied by additional easy-axis contribution along the high-symmetry tetragonal axis and in a certain range of biquadratic exchange constant, the phase diagrams similar to the experimental one of GdRu2Si2 [37, 21, 31] can be obtained. Thus, we also obtain a reasonable estimation for the biquadratic exchange strength.
@@
The rest of the paper is organized as follows. In Section II, we introduce the formalism. Various possible spin structures and relations between their free energies are discussed in Sec. III. In Sec. IV, we present particular phase diagrams, including the one corresponding to GdRu2Si2. Section V is devoted to a general discussion of the results. Section VI contains our conclusions and summary. Finally, two appendices contain detailed information about free energy calculations and topological properties of some spin structures.
II General formulae
@@
Following the previous studies [25, 26, 27], we consider a model which includes frustrated isotropic exchange, magnetodipolar interaction and/or anisotropic exchange, Zeeman term, and biquadratic exchange. Assuming one magnetic ion in the unit cell, we can write the reciprocal space Hamiltonian as follows:
@@
where $N$ is the number of lattice sites and we measure magnetic fields in energy units, $h=g\mu_{\textrm{B}}H$. The last term with $K>0$ here is written in the form suggested in Refs. [30, 25] as the most important contribution to the biquadratic exchange in itinerant magnets. The first two terms (3) and (4) can be combined into a single term $\mathcal{H}_{\textrm{EX}}+\mathcal{H}_{\textrm{AN}}=-\sum_{\mathbf{q}}\mathcal{H}^{\alpha\beta}_{\mathbf{q}}S^{\alpha}_{\mathbf{q}}S^{\beta}_{-\mathbf{q}}$. Noteworthy, tensor $\mathcal{H}^{\alpha\beta}_{\mathbf{q}}$ eigenvalues $\lambda_{1}(\mathbf{q})\geq\lambda_{2}(\mathbf{q})\geq\lambda_{3}(\mathbf{q})$ and respective eigenvectors $\mathbf{v}_{1}(\mathbf{q}),\,\mathbf{v}_{2}(\mathbf{q}),\,\mathbf{v}_{3}(\mathbf{q})$ define the “zoo” of possible spin structures in each particular model.
@@
We aim to discuss experimentally relevant cases, when the possible modulation vectors lie in-plane [we denote them as $\mathbf{k}_{x}=(k,0,0)$ and $\mathbf{k}_{y}=(0,k,0)$]. They are equivalent due to tetragonal symmetry [21, 31]. These vectors are roughly (up to small anisotropy-induced corrections) determined by the maxima of the exchange interaction Fourier transform. In the discussion below, we will also need $\mathbf{k}_{\pm}=\mathbf{k}_{x}\pm\mathbf{k}_{y}$. In GdRu2Si2, $k=0.22\times 2\pi/a$, where $a$ is the in-plane lattice parameter [21]. Moreover, the corresponding spin structures have transverse order parameters, and we conclude that the longitudinal directions for $\mathbf{k}_{x}$ and $\mathbf{k}_{y}$ are the respective hard axes. Then, the corresponding perpendicular in-plane directions can be easy or middle axes. Importantly, the former case is relevant to magnetodipolar interaction, which should be important for Gd3+ ions with zero orbital momentum [26]. In the general case, the anisotropic part of the Hamiltonian can also include compass anisotropic exchange [27] and single-ion terms. Fig. 1 illustrates these crucial concepts.
Figure 2: Sketch of main spin structures in the xyxy plane, which can appear in typical phase diagrams of tetragonal frustrated antiferromagnets. Color encodes $\hat{z}$ projections of spins (red – “spin up”, violet – “spin down”, green – spin in the xyxy plane). (a) Superposition of two spin-density waves with modulation vectors $\mathbf{k}_{x}$ and $\mathbf{k}_{y}$ and perpendicular in-plane polarizations (2SXY). In real space, it corresponds to a vortex lattice. (b) Superpositions of two SDWs with in-plane and out-of-plane polarizations (S2YZ). (c) Superposition of screw helicoid with $\mathbf{k}_{y}$ and SDW with $\mathbf{k}_{x}$ (1Q+1S); meron – anti-meron lattice. (d) Topologically-nontrivial square skyrmion lattice made of two screw helicoids with $\mathbf{k}_{x}$ and $\mathbf{k}_{y}$, harmonics with $\mathbf{k}_{\pm}$ polarized along $\hat{z}$, and uniform magnetization (2Q). Its topological charge per unit cell is $n_{\textrm{sk}}=1$ (e) Topologically-trivial version of 2Q structure at stronger external field (all $s^{z}_{i}>0$), $n_{\textrm{sk}}=0$. (f) Alternative topologically-nontrivial structure 2Q with $n_{\textrm{sk}}=2$.
@@
For our calculations, we employ the mean-field approach based on the smallness of the thermodynamically average ordered spin value, $s=\langle S\rangle\ll S$. In this case, which corresponds to high temperatures near $T_{\textrm{N}}$, we can express the free energy as a function of $s_{\mathbf{q}}$ (reciprocal space) and $\mathbf{s}_{i}$ (direct space) as follows (see, e.g., Ref. [38] for details):
$\mathcal{F}$ $=$ $-\sum_{\mathbf{q}}\mathcal{H}^{\alpha\beta}_{\mathbf{q}}s^{\alpha}_{\mathbf{q}}s^{\beta}_{-\mathbf{q}}-\sqrt{N}\mathbf{h}\cdot\mathbf{s}_{\bf 0}+AT\sum_{i}s^{2}_{i}$ (7)
$+\frac{K}{N}\sum_{\mathbf{q}}(\mathbf{s}_{\mathbf{q}}\cdot\mathbf{s}_{-\mathbf{q}})^{2}+BT_{\textrm{N}}\sum_{i}s^{4}_{i},$
@@
For $S=7/2$, one has $A\approx 0.095$ and $B\approx 0.002$. Note that by assuming a certain spin structure, we readily obtain its free energy (7) in the form of a quartic polynomial suitable for either analytical or numerical minimization [38, 26, 35].
III Spin structures and phase transitions
In the present study, we consider the magnetic field only along the $\mathbf{c}$ axis. Some of the relevant spin structures are shown in Fig. 2. We distinguish: (i) single-modulated spin-density wave (SDW) which we denote 1S, it turns out to be relevant only for easy axes along $\hat{z}$, (ii) three variants of double-modulated superpositions of two SDWs with different polarizations (2SXY, 2SYZ, 2SZZ), (iii) vertical screw helicoid (1Q), (iv) transverse cone spiral (TC), (v) hybrid structure made of vertical spiral and SDW with perpendicular modulation vector (1Q+1S), (vi) superposition of two vertical spirals (2Q), which is the square skyrmion lattice in a certain range of fields, (vii) alternative 2Q ordering (2Q), where $\hat{z}$ components of the spin structure have $\mathbf{k}_{\pm}$ modulation vectors. Details of calculations supporting our claims are presented in Appendix A.
@@ III Spin structures and phase transitions
Moreover, it can be shown that in the region where 1Q+1S does not exist $t<(5b+K)\Lambda_{1}/2b$], but 1Q does $t>(3b+K)\Lambda_{1}/2b$], 2SXY has minimal free energy. As a corollary, the 1Q structure does not appear in the phase diagrams.
The same applies if we compare the free energies of 1Q+1S (51) and 2Q (55) [without the correction (59) important at moderate magnetic fields]. Note that these phases are defined in the same region $t>(5b+K)\Lambda_{1}/2b$ and emerge via the second-order phase transition from 2SXY. The difference in the free energies
$f_{\textrm{2Q}}-f_{\textrm{1Q+1S}}=\frac{(2bt-5b\Lambda_{1}-K\Lambda_{1})^{2}}{8(9b+2K)(2b^{2}+5bK+K^{2})}\geq 0.$
@@
$\mathbf{s}_{j}$ $=$ $s(\hat{y}\sin{\mathbf{k}_{x}\mathbf{R}_{j}}+\hat{x}\sin{\mathbf{k}_{y}\mathbf{R}_{j}})$ (12)
$+\hat{z}[p(\cos{\mathbf{k}_{x}\mathbf{R}_{j}}+\cos{\mathbf{k}_{y}\mathbf{R}_{j}})+m],$
with certain $s,p$ and $m$ (see Appendix A). $z$-components of spin structure due to the quartic term in the free energy (7) lead to instability for the emergence of additional harmonics $\mathbf{k}_{\pm}$ lowering the total free energy. The physics here resembles the conventional mechanism of the stabilization of hexagonal skyrmion lattices in magnetic field due to the interplay of constant magnetization and three helical constituents with $\mathbf{k}_{1}+\mathbf{k}_{2}+\mathbf{k}_{3}=0$ (see, e.g., Refs. [38, 7, 18, 35]). In the present case, the crucial terms $\propto mp^{2}$ emerge due to the equality $\mathbf{k}_{x}\pm\mathbf{k}_{y}-\mathbf{k}_{\pm}=0$ and similar ones. Notably, the corresponding correction to the spin structure
@@
“patches” the holes (where $\mathbf{s}_{j}=m\hat{z}$) in the modulated part of (12) for $\mathbf{k}_{x,y}\mathbf{R}_{j}=\pi l_{x,y}$ and $\mathbf{k}_{x}\mathbf{R}_{j}-\mathbf{k}_{y}\mathbf{R}_{j}=2\pi l+\pi$ (all $l$s are integers). Importance of these harmonics was also highlighted in Refs. [39, 36].
In the external field, the free energy parameters should be altered as follows: $t\rightarrow t^{\prime}=t-2b\chi^{2}h^{2}$ and $\Lambda_{1}\rightarrow\Lambda^{\prime}_{1}=\Lambda_{1}+4b\chi^{2}h^{2}$, where $\chi$ represents the uniform susceptibility. All free energies also acquire a “paramagnetic” correction $\chi h^{2}/2$. So, in the external field, the above statements are intact. However, note that the correction to the 2Q phase free energy (59) is $\propto m^{2}$ and becomes crucial at moderate fields for which both $m$ and $p$ are significant.
@@
Figure 3: Phase diagrams for simplified models, i.e., for which we include only one of the crucial ingredients. (a) Without the biquadratic exchange, but including anisotropy due to magnetodipolar interaction, the skyrmionic 2Q phase is stable only in a tiny region. At a lower $T$, it loses competition to the transverse conical spiral. (b) For nonzero biquadratic exchange with $K=0.05$ K but without anisotropy, TC is stable in a large region of moderate fields. (c) The same as (b), but $K=0.1$ K. Larger $K$ allows substituting TC with vortical phase 2SXY and makes the square skyrmion lattice stable in the intermediate field region.
Importantly, near the transition point from 2SXY, for 2Q structure $p\ll 1$ (and, hence, $s_{\pm}\ll 1$), uniform magnetization $m$ dominates in $s^{z}_{i}$ and 2Q structure is topologically trivial, see Appendix B for details. The condition of nonzero topological charge is $2(p+s_{\pm})>m$. Note also that the transverse cone spiral, involving the hard magnetization axis, can be stabilized at lower $T$ [it is governed by the $\Lambda_{2}>\Lambda_{1}$ parameter, see Eq. (47)]. Its competition with the other phases should be checked separately. Particular examples of phase diagrams are presented below.
@@
IV Phase diagrams and application to $\text{Gd}\text{Ru}_{2}\text{Si}_{2}$